OpenStudy (yacoub1993):

Find the roots of the polynomial equation. -x^3 + 5x^2 - 11x + 55 = 0

OpenStudy (yacoub1993):

@gorv

OpenStudy (yacoub1993):

http://prntscr.com/33o12o A am i right

OpenStudy (gorv):

just take - x^2 common in first 2 terms and -11 in last 2

OpenStudy (yacoub1993):

A or B not sure

OpenStudy (gorv):

-(x^3-5x^2+11x-55)=0 -(x^2(x-5)+1(x-5))=0 x=5 and x^2=-11 D

OpenStudy (gorv):

@Yacoub1993 got??

OpenStudy (yacoub1993):

Yes i got it ..... i miscalculated it so i got -5 for x

OpenStudy (yacoub1993):

I got it

OpenStudy (gorv):

okk